OpenStudy (anonymous):

Twenty-three times a positive integer is more than its square by 63. Find the integer.

OpenStudy (anonymous):

you can say like that 23x - x^2 = 63

OpenStudy (anonymous):

welcome to openstudy

OpenStudy (waleed_imtiaz):

23x=63+x^2 x^2-23x+63=0 solve for x........

OpenStudy (anonymous):

you have calculate the quadratic equation

OpenStudy (anonymous):

@DKS are u here